# Know Your Neighborhood: General and Zero-Shot Capable Binary Function Search Powered by Call Graphlets

### v1

@@ 3.3.3 Data Preparation
A uniform data preparation pipeline was implemented to process all datasets outlined above. This process can be broken down into five key stages: extraction, fusion, deduplication, augmentation and sub-setting.
Extraction: The initial stage involved extracting call graphs and function metadata for each binary within the datasets. This was achieved by utilizing the bin2ml tool which uses radare2 as the underlying software reverse engineering tool to load, process and output the required data.
Fusion: Following extraction, the extracted call graph and function metadata were combined. This process resulted in the creation of function-level call graphlets. Each call graphlet encapsulated the functions local call graph structure whereby each node’s features was it’s corresponding function’s metadata. Each call graphlet was output as a Networkx [7] compatible JSON object to support seamless loading into our chosen graph neural network framework, PyTorch Geometric [5].
@@
    Call graphlets are initially grouped based on the software binary from which they originated. This binary-specific grouping facilitates a more efficient deduplication process.
@@
    Subsequently, each call graphlet within a group is hashed using a suitable hashing algorithm. The resulting hash values are then stored within a hash set. This hash set serves as a mechanism to track unique call graphlets encountered for each binary.
@@
All of the remaining, unique call graphlets are saved and form the basis of the given dataset used to train or evaluate our proposed approach.
Augmentation: Once deduplication was complete, each graph was loaded into Networkx and the edge betweenness values were generated and then appended to the graph as floating point edge weights.
Subsetting: The last stage focused on sub-setting the data. This was not necessary for all datasets but in the case of BinkitNormal and BinkitNoInline, due to the number of unique functions call graphlets generated, it was not feasible to evaluate a model on all functions present due to hardware limitations. For both, we used a random sampling strategy to create test datasets. Each sample contained a total of 1 million functions. This subset was created once for each source dataset and then saved for re-use. Both of these splits are provided as part of the release of this research.
@@ 3.4 Model
The model used within our approach is a bespoke 3-layer graph neural network incorporating several different GNN components to create a novel GNN architecture. A diagram visually representing the model can be seen in Figure 2
GraphConv Message Passing GNN
The graph convolution operator used within this study is the operator proposed in [16]. It computes the updated feature representation $x_{i}^{\prime}$ for a node $i$ in a graph by combining two components: 1) a linear transformation of the node’s own input features $W_{1}x_{i}$, and 2) a weighted sum of the features of its neighboring nodes $\sum_{j\in N(i)}e_{j,i}\cdot x_{j}$, where the neighbors’ features are first weighted by the strength of their connections $e_{j,i}$ and then linearly transformed $W_{2}$. The resulting output feature vector $x_{i}^{\prime}$ captures both the node’s intrinsic features and the information from its local neighborhood. This operator has been shown to be able to capture higher-order graph structures and outperform a range of standard GNN and graph kernel approaches. This is due to approach considering structures within the graphs at varying scales.
@@
Instance Normalization
We adopt Instance Normalization (IN) [22] as the graph normalization technique omitting the learnable scale and shift parameters $\gamma$ and $\beta$, as described in Equation 3. IN is applied to the node features, where $x_{i}\prime$ represents the normalized node feature for node $i$. It centers the features by subtracting the mean $E[x]$ across all nodes within a given graph, scales them by dividing by the square root of the variance $Var[x]$ plus a small constant $\epsilon$ (set to $051e-$ 05). This normalization technique aims to stabilize the training process by reducing the internal covariate shift, ensuring consistent distribution of node features across the layers of the graph neural network.
$x_{i}{\prime}=\frac{x-E[x]}{\sqrt{Var[x]+\epsilon}}$ (3)
Due to the cross-architecture context of our research, the selection of architecture, compiler, and optimization combinations can significantly influence a function’s structure and form. IN exhibits an interesting property wherein it performs normalization on each function independently at the graph level. This means that variations among features within our examples are handled separately, which is crucial in our domain due to substantial variations across multiple features, such as the number of instructions, total number of edges, in-degree, and out-degree.
Softmax Aggregation
Similar to other recent works such as HermesSim [8], we opt use the Softmax aggregation function as defined as follows:
$softmax(X|t)=\sum_{x_{i}\in X}\frac{exp(t\cdot x_{i})}{\sum_{x_{j}\in X}exp(t\cdot x_{j})}\cdot x_{i}$ (4)
where $t$ controls the softness of the softmax aggregation over the input feature set $X$, which in our case is set to 1. This differs from the usage within HermesSim in two key ways. Firstly, we adopt the vanilla version of softmax aggregation which omits the multi-headed attention extensions. And secondly, we use the softmax aggregation function as the GNN’s aggregation function as opposed to using it as a global graph pooling operator. The softmax operator is used to create node-level representations at each message passing step.
Layer-Wise Connections + Max Pool
We opt to have Layer-Wise feature connections from each of the graph neural network layers to store the node level representations from each convolution. The node-level representations are stored from each layer-wise connection before each of them being max pooled to create a graph level representation where the resulting feature vector is a feature-wise maximum across the node features of that layer. This operation can be seen in Equation 5.
@@
Each of the max pooled graph-level embeddings are then concatenated together $X_{c}=(X_{1},X_{2},X_{3})$ resulting in a single vector of size $3*X_{input}$. This feature vector is then passed into a two layer linear network, mapping from $3*hidden_{d}\to hidden_{d}\to\frac{hidden_{d}}{2}$. The output of this is then normalized using $F2$ normalization to produce the final embedding.
@@ 3.5 Training
The model was trained using pair-based deep metric learning using a BatchHard mining scheme [9]. We paid special attention to the sampling methodology to ensure that each batch contained multiple (2 in our case) versions of any function present. This ensured that it was always possible to create a positive pair for each function present as well as provide a rich range of possible negative samples.
The model’s configuration was a hidden dimension size of 256 and an output embedding dimension of 128. Circle loss [21] was used as the loss function configured with $m$ set to 0.25 and $\gamma$ set to 256. Circle loss and the hyper parameters were chosen due to good results presented in previous works [4]. The model was optimized using the Adam optimizer [12]. The model was trained for 350 epochs with each epoch having 100K graphs randomly sampled from the total training dataset. A cosine simulated annealing with warm restarts learning rate schedule starting at 0.0005 and finishing at 0.0001 was used and each batch had 256 graphs within it. The number of epochs before warm restart started at 50 and doubled for 2 subsequent times (50 -> 100 -> 200) for a total of 350 training epochs.
@@ 4 Evaluation
We implement KYN using PyTorch Geometric [5]. We use bin2ml alongside radare2 to generate the data for all the experiments outlined below. All experimentation was conducted on a workstation with 64GB RAM alongside a NVIDIA 4090 24GB GPU.
@@ 4.1 Tasks
For both cross and mono-architecture evaluation, we provide results across a range of search pool sizes ranging from 100 to 10000. The search pool size of 100 and 10000 are used to compare against other SOTA approaches with the remaining provided for future researchers to compare against.
@@
FASER NRM [4] ESIL IR Strings 0.51 0.57 - -
@@
KYN (Ours) Call Graphlets + Metadata 0.76 0.80 0.45 0.50
Table 2: Dataset-1 (Cisco Talos Binary Similarity [15]) Test Results across search pools of 100 and 10,000 compared against relevant benchmarks
@@ 4.2 Metrics
To facilitate robust comparison, we adopt the same metrics used within the domain of Recall at 1 (R@1) and Mean Reciprocal Rank at 10 (MRR@10) [15]. These metrics are used for all results presented as part of RQ1, RQ2 and RQ3. At points throughout reporting the results for RQ1, RQ2 and RQ3 we also report Normalized Discount Cumulative Gain at 10 (NDCG@10). NDCG is an alternative metric to MRR and unlike MRR, considered results throughout the K selected rather than just the first occurrence. The NDCG@10 scores have been provided to aid future researchers compare against our approach. For RQ4, we report the ranks at which the vulnerable function was present at after the search was conducted. This is the same as previous studies [15]. We also report the mean and median rank to support comparison in a similar manner to [4].
@@ 4.3 Baseline Approaches
    GMN [14] is a variant of the GNN proposed in the same paper and have been shown to outperform alternative baseline graph neural network approaches [15]. This approach does however rely on 1-to-1 matching across the entire comparison corpus which is a significant drawback.
@@
    HermesSim [8] is a state of the art approach that leverages a novel graph data structure named Semantic Orientated Graph alongside a graph neural network to embed functions before then comparing them.
For the mono-architecture experiments, we also compare against jTrans [23]. jTrans is a transformer model trained using a novel jump orientated pre-training task on a large x86-64 dataset and can be considered as one of the state of the art models for x86-64 mono-architecture binary function search.
@@ 4.4.1 Dataset 1
Trex 512 Tokens 0.46 0.53
GNN (from athors of GMN) CFG + BoW opc 200 0.57 0.67
GMN CFG + BoW opc 200 0.61 0.71
KYN (Ours) Call Graphlets 0.76 0.82
Table 3: Dataset-2 Test Results across search pools of 100 compared against relevant benchmarks. Comparison scores taken directly from [15]
The results presented in the Table 2 show the results generated for search pools 100 and 10000. Starting first with the results for the smaller, 100 negative search pool size, our approaches performs comparably to the current state of the art approaches, HermesSim [8] and BinFinder [19] with reported metrics that are near identical. When comparing against our other chosen SOTA GNN baselines, GraphMoco [20] and GMN [14], our approach significantly outperforms across all collected metrics with a R@1 relative increase of 32% and 51.2% and a MRR@10 relative increase of 28.6% and 38.8% respectively. When compared against two state of the art cross-architecture Transformer models FASER [4] and Trex [18], our approach again significantly outperforms them with a R@1 relative increase of 39.8% and 104% and a MRR@10 relative increase of 33.6% and 80.7%. Turning now to the larger search pool size of 10000, our proposed approach again reports near identical results to HermesSim with a marginally better R@1 but a lower MRR@10 suggesting both approaches are performing similarly.
In addition to the search pools sizes of 100 and 10,000, we also present results for the search pool sizes of 250 and 1000 in Table 3. This was primarily due to how significant the difference is for the results between 100 and 10,000 is and provides a means of reasoning about the performance degradation as the pool size increases as well as providing future researchers empirical results to compare against.
The results presented in Table 3 show that, as expected, all metrics degrade as the search pool size increases, with the degradation occurring steadily. Notably however, our approach even outperforms the bottom four baselines (Trex, GMN, FASER, GraphMoco) when faced with search pools containing 1,000 functions compared to when each of the approaches is dealing with only 100 functions. This demonstrates the effectiveness of our approach, especially when faced with a larger number of possible candidate functions to search.
@@ 4.4.2 Dataset-2
The results shown Table 4 demonstrate that our approach significantly outperforms all the baseline approaches. Compared to the best baseline, our method achieves a 22.9% increase in R@1 (recall at 1) and a 14.8% increase in MRR@10 (mean reciprocal rank at 10). Interestingly, the performance of our approach is comparable to the results we obtained for Dataset-1. This is not the case for the baseline approaches, which perform better on this dataset than on Dataset-1. This suggests that not only does our model outperform the baselines in terms of the metric scores, but it also exhibits more consistent performance across different datasets.
@@
250 0.670 0.711 0.735
1000 0.572 0.616 0.641
Table 4: Results for Dataset-1 Across Different Search Pools
In a similar manner to Dataset-1, we also generated results for search pools of 250, 1000 and 10,000. The results for this are presented in Table 5. As the search pool size increases, the performance degradation is more significant. From theorising why this may be, we think this is likely due to the composition of the Dataset-1 test set. The test set split is dominated by the Z3 (in terms of on-disk size and number of functions) resulting in an intrinsically varied sample. When faced with a dataset with more scope for similar functions from different binaries and when at large pool sizes, the task becomes more challenging therefore reducing overall performance.
@@
250 0.674 0.737 0.765
1000 0.566 0.641 0.679
10000 0.328 0.417 0.461
Table 5: Results for Dataset-2 Across Different Search Pools
@@ 4.4.3 Binkit Noinline
The results from evaluating our approach for binkit-noinline architecture presented in Table 6. Our approach again performs consistently across all of the metrics collected and reports similar results to those collected when using Dataset-1 and Dataset-2.
@@
100 0.824 0.873 0.893
250 0.774 0.831 0.858
1000 0.636 0.714 0.751
10000 0.436 0.529 0.574
Table 6: Binkit-Noinline Metrics Across Different Search Pools
@@ 4.4.4 Summary
In summary, the results above demonstrate that, after robust evaluation across a range of different binary benchmark datasets, that our approach performs equally or significantly better to comparison benchmarks. The breadth of datasets used for evaluation demonstrates that our approach could be considered a general approach which can generalize to unseen binaries within a cross-architecture setting.

### v2

@@ 3.3.3 Data Preparation
A uniform data preparation pipeline was implemented to process all datasets outlined above except the vulnerability benchmark. This process can be broken down into five key stages: extraction, fusion, deduplication, augmentation and sampling.
Extraction: The initial stage involved extracting call graphs and function metadata for each binary within the datasets. This was achieved by utilizing the bin2ml[13] tool which uses radare2[16] as the underlying software reverse engineering tool to load, process and output the required data.
Fusion: Following extraction, the extracted call graph and function metadata were combined. This process resulted in the creation of function-level call graphlets. Each call graphlet encapsulated the functions local call graph structure whereby each node’s features was it’s corresponding function’s metadata. Each call graphlet was output as a Networkx [17] compatible JSON object to support loading into our chosen graph neural network framework, PyTorch Geometric [18].
@@
    Initially call graphlets are grouped based on the software binary from which they originated. This binary-specific grouping facilitates a more efficient deduplication process.
@@
    Subsequently, each call graphlet within a group is hashed. The resulting hashes are then stored within a hash set. This hash set serves as a mechanism to track unique call graphlets encountered for each binary.
@@
All the remaining unique call graphlets are saved and form the basis of the given dataset used to train or evaluate our proposed approach.
Augmentation: Once deduplication was complete, each remaining graph was loaded into Networkx and the edge-betweenness centrality calculated and then appended to the graph as an edge weight.
Sampling: The last stage focused on sampling the data. This was only necessary for BinkitNormal and BinkitNoInline. Due to the scale of the datastets once processed, it was not feasible to evaluate the model of all functions due to hardware limitations. To overcome this, we used a random sampling strategy to create test datasets. Each sample contained a total of 1 million functions. This subset was created once for each source dataset and then saved for re-use. Both of these splits are provided as part of the release of this research.
@@ 3.4 Model
The model used within our approach is a 3-layer graph neural network incorporating several GNN components to create a novel GNN architecture. A diagram visually representing the model can be seen in Figure 2
GraphConv Message Passing GNN: The graph convolution operator used within this study is the operator proposed in [19] and is shown in Equation 2. It computes the updated feature representation $x_{i}^{\prime}$ for a node $i$ in a graph by combining two components: 1) a linear transformation of the node’s own input features $W_{1}x_{i}$, and 2) a weighted sum of the features of its neighboring nodes $\sum_{j\in N(i)}e_{j,i}\cdot x_{j}$, where the neighbors’ features are first weighted by the strength of their connections $e_{j,i}$ and then linearly transformed by $W_{2}$. The resulting output feature vector $x_{i}^{\prime}$ captures both the node’s intrinsic features and the information from its local neighborhood. This operator has been shown to be able to capture higher-order graph structures and outperform a range of standard GNN and graph kernel approaches. This is due to approach considering structures within the graphs at varying scales. The properties of this operator suit our call graphlet data representation and help leverage the richness of a functions calling relationships.
@@
Instance Normalization: We adopt Instance Normalization (IN) [20] as the graph normalization technique omitting the learnable scale and shift parameters $\gamma$ and $\beta$, as described in Equation 3. IN is applied to the node features, where $x_{i}\prime$ represents the normalized node feature for node $i$. It centers the features by subtracting the mean $E[x]$ across all nodes within a given graph, scales them by dividing by the square root of the variance $var[x]$ plus a small constant $\epsilon$ (set to $051e-$ 05). This normalization technique aims to stabilize the training process by reducing the internal covariate shift, ensuring consistent distribution of node features across the layers of the graph neural network.
[…]
The selection of architecture, compiler, and optimization combinations can significantly influence a function’s structure and form therefore introducing a large variance in our simple, node level statistical features. IN exhibits an interesting property wherein it performs normalization on each function independently at the graph level. This means that variations among features within our examples are handled separately, which is crucial due to substantial variations across multiple features, such as the number of instructions, total number of edges, in-degree, and out-degree.
Softmax Aggregation: Similar to other recent works such as HermesSim [10], we use the Softmax aggregation function as defined as follows:
[…]
Where $t$ controls the softness of the softmax aggregation over the input feature set $X$, which we set to 1. This differs from the usage within [10] in two key ways. Firstly, we adopt the vanilla version of softmax aggregation which omits the multi-headed attention extensions introduced in [10]. And secondly, we use the softmax aggregation function as the GNN’s aggregation function as opposed to using it as a global graph pooling operator. The softmax operator is used to create node-level representations at each message passing step.
Layer-Wise Connections + Max Pool: We opt to have Layer-Wise feature connections from each of the model layers to store the node level representations from each convolution. The node-level representations are stored from each layer-wise connection before each of them being max pooled to create a graph level representation where the resulting feature vector is a feature-wise maximum across the node features of that layer. This operation can be seen in Equation 3.
@@
The max-pooled graph-level embeddings are concatenated as:
$X_{c}=[X_{1}\parallel X_{2}\parallel X_{3}]\in\mathbb{R}^{3d_{in}}$
Where $\parallel$ denotes concatenation and $d_{in}$ is the intermediate hidden dimension. This feature vector is then transformed through a two-layer linear network:
$h_{1}=f_{1}(X_{c})\in\mathbb{R}^{3d_{in}}$
$h_{2}=f_{2}(h_{1})\in\mathbb{R}^{3d_{in}/2}$
The final embedding is obtained through L2 normalization:
$z=\frac{h_{2}}{|h_{2}|_{2}}$
@@ 3.5 Training
The model was trained using pair-based deep-metric learning alongside a BatchHard mining scheme [21]. We paid special attention to the sampling methodology to ensure that each batch contained multiple (2 in our case) versions of any function present. This ensured that it was always possible to create a positive pair for each function present as well as provide a rich range of possible negative samples.
The model’s configuration was a hidden dimension size of 256 and an output embedding dimension of 128. Circle loss [22] was used as the loss function configured with $m$ set to 0.25 and $\gamma$ set to 256. Circle loss and the hyperparameters were chosen due to good results presented in previous works [5]. The model was optimized using the Adam optimizer [23]. The model was trained for 350 epochs with each epoch having 100K graphs randomly sampled from the total training dataset. A cosine simulated annealing with warm restarts learning rate schedule starting at 0.0005 and finishing at 0.0001 was used, and each batch had 256 graphs within it. The number of epochs before warm restart started at 50 and doubled for 2 subsequent times (50 -¿ 100 -¿ 200) for a total of 350 training epochs.
@@ 4 Evaluation
We implement KYN using PyTorch Geometric [18]. We use bin2ml[13] alongside radare2[16] to generate the data for all the experiments outlined below. All experimentation was conducted on a workstation with 64GB RAM alongside a NVIDIA 4090 24GB GPU.
@@ 4.1 Tasks
For both cross and mono-architecture evaluation, we provide results across a range of search pool sizes ranging from 100 to 10,000. The search pool size of 100 and 10,000 are used to compare against other SOTA approaches with the remaining provided for future researchers to compare against.
@@
FASER NRM [5] ESIL IR Strings 0.51 0.5 - -
@@
KYN (Ours) Call Graphlets + Metadata 0.75 0.79 0.45 0.50
TABLE II: Dataset-1 (Cisco Talos Binary Similarity [12]) Test Results across search pools of 100 and 10,000 compared against relevant benchmarks
@@ 4.2 Metrics
To facilitate robust comparison, we adopt the same metrics used within the domain of Recall at 1 (R@1) and Mean Reciprocal Rank at 10 (MRR@10) [12]. These metrics are used for all results presented as part of RQ1, RQ2 and RQ3. At points throughout reporting the results for RQ1, RQ2 and RQ3 we also report Normalized Discount Cumulative Gain at 10 (NDCG@10). The NDCG@10 scores have been provided to aid future researchers compare against our approach. For RQ4, we report the ranks at which the target vulnerable function was present at after the search was conducted. This is the same as previous studies [12]. We also report the mean and median rank to support comparison similarly to Collyer[5].
@@ 4.3 Baseline Approaches
    GMN [24] is a one of the GNN methods proposed in [24] and has been shown to outperform alternative baseline graph neural network approaches [12]. This approach does however rely on 1-to-1 matching across the entire comparison corpus which is a significant drawback.
@@
    HermesSim [10] is a state-of-the-art approach that leverages a novel graph data structure named Semantic Orientated Graph alongside a graph neural network to embed functions before then comparing them.
For the mono-architecture experiments, we also compare against jTrans [4]. jTrans is a transformer model trained using a novel jump- orientated pre-training task on a large x86-64 dataset and can be considered as one of the state-of-the-art models for x86-64 mono-architecture binary function search.
@@ 4.4.1 Dataset 1
Trex [6] 512 Tokens 0.46 0.53
GNN [24] CFG + BoW opc 200 0.57 0.67
GMN [24] CFG + BoW opc 200 0.61 0.71
KYN (Ours) Call Graphlets 0.75 0.81
TABLE III: Dataset-2 Test Results across search pools of 100 compared against relevant benchmarks. Comparison scores taken directly from [12]
The results presented in the Table II show the results generated for search pools 100 and 10,000. Starting first with the results for the smaller, 100 search pool size, our approach performs comparably to the current state-of-the-art approaches, HermesSim [10] and BinFinder [11] with reported metrics that are near identical. When comparing against our other chosen SOTA GNN baselines, GraphMoco [25] and GMN [24], our approach significantly outperform them across all collected metrics with a R@1 relative increase of 27% and 42% and a MRR@10 relative increase of 24% and 32% respectively. When compared against two state of the art cross-architecture Transformer models FASER [5] and Trex [6], our approach again significantly outperforms them with a R@1 relative increase of 32% and 68% and a MRR@10 relative increase of 37% and 57%. Turning now to the larger search pool size of 10,000, our proposed approach again reports near identical results to HermesSim with a marginally better R@1 but a lower MRR@10 suggesting both approaches are performing similarly.
In addition to the search pools sizes of 100 and 10,000, we also present results for the search pool sizes of 250 and 1,000 in Table IV. This was primarily due to how significant the difference is for the results between 100 and 10,000 is and provides a means of reasoning about the performance degradation as the pool size increases as well as providing future researchers empirical results to compare against.
The results presented in Table IV show that, as expected, all metrics degrade as the search pool size increases, with the degradation occurring steadily. Notably however, our approach outperforms the bottom four baselines (Trex, GMN, FASER, GraphMoco) when faced with search pools containing 1,000 functions compared to when each of the approaches is dealing with only 100 functions. This demonstrates the effectiveness of our approach, especially when faced with a larger number of possible candidate functions to search.
@@ 4.4.2 Dataset-2
The results shown in Table III demonstrate that our approach significantly outperforms all the baseline approaches on Dataset-2. Compared to the best baseline, our method achieves a 19% increase in R@1 and a 12% increase in MRR@10. Interestingly, the performance of our approach is comparable to the results we report for Dataset-1. This is not the case for the baseline approaches, which perform better on this dataset than on Dataset-1. This suggests that not only does our model outperform the baselines in terms of raw scores, but it also exhibits more consistent performance across different datasets which suggests enhanced generalizability.
@@
250 0.664 0.703 0.726
1,000 0.554 0.611 0.638
TABLE IV: Results for Dataset-1 Across Different Search Pools
Similarly to Dataset-1, we also generated results for search pools of 250 and 1,000. The results for this are presented in Table V. As the search pool size increases, the performance degradation is more significant. We think this is likely due to the composition of the Dataset-1 test set. The test set split is dominated by the Z3 library (in terms of on-disk size and number of functions) resulting in a test sample with limited variety. When faced with a dataset with more scope for similar functions from different binaries and when at large pool sizes, the task becomes more challenging therefore reducing overall performance.
@@
250 0.639 0.716 0.753
1,000 0.536 0.619 0.660
TABLE V: Results for Dataset-2 Across Different Search Pools
@@ 4.4.3 Binkit Noinline
The results from evaluating our approach for binkit-noinline architecture presented in Table VI. Our approach again performs consistently across all the metrics collected and reports similar results to those collected when using Dataset-1 and Dataset-2.
@@
100 0.828 0.874 0.895
250 0.777 0.834 0.860
1,000 0.663 0.728 0.759
10,000 0.442 0.534 0.582
TABLE VI: Binkit-Noinline Metrics Across Different Search Pools
@@ 4.4.4 Summary
The results above answer RQ1 and demonstrate that, after robust evaluation across a range of different binary function search benchmark datasets, that our approach performs equally or significantly better when compared against baseline approaches. The breadth of datasets used for evaluation also demonstrates that our approach could be considered a general approach which can generalize to unseen binaries, compiler options and compilers.